Miguel is 44 years older than Kendra. In 14 years he will be 2 times her age. How old is each now?
Answer by ikleyn(52781)

.
Miguel is 44 years older than Kendra. In 14 years he will be 2 times her age. How old is each now?
~~~~~~~~~~~~~~~~~~~~~~

M = K + 44,         (1)   ("Miguel is 44 years older than Kendra")
M + 14 = 2(K+14).   (2)   ("In 14 years he will be 2 times her age.)

Substitute (1) into (2). You will get

(K+44) + 14 = 2(K+14).

Simplify and solve:

K + 58 = 2K + 28  --->  K = 30.

Answer.  Kendra is 30 years old. Miguel is 74.

There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - A fresh formulation of a traditional age problem
    - Really intricate age word problem
    - Selected age word problems from the archive
in this site.

Read them and become an expert in solving age problems.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

It contains a lot of other word problems and many other interesting and useful things.